# references.bib

### v2

@@ 4 Conclusion and discussion
Generally, performance increases of GPT-4o and similar models are to be expected in the future, further amplifying the viability of this approach. To reduce dependency on OpenAI and monetary cost, future work should evaluate open-source alternatives. Similarity search is more efficient in computational and monetary terms, therefore currently remaining a viable alternative despite inferior results. The results varied between classes and showed the worst performance for paving stones - bad, where there was no improvement to the baseline. Further experiments, especially with more images for the smallest class, as well as advancements, could be explored, e.g., by incorporating different clustering strategies [vo_automatic_2024].
We envision extensions of the dataset to include road-type information as well as unsuitable images, such as blurred ones or those without a road in focus. Including these will enable training models to automatically check image suitability, thus improving classification robustness and enhancing the GPT-4o prompting strategy.
[…]
Appendix A Labeling scheme
A.1 Asphalt
Quality Description
excellent As good as new asphalt, on which a skateboard or rollerblades will have no problem.
good Asphalt showing the first signs of wear, such as narrow, smaller than 1.5 cm cracks, or wider cracks filled up with tar, shallow dents in which rainwater may collect, which may cause trouble for rollerblades but not for racing bikes.
intermediate Asphalt roads that show signs of maintenance, such as patches of repaired surface, wider cracks larger than 2cm. Asphalt sidewalks may contain potholes, but these are small, shallow (<3cm deep) and can be easily avoided. Asphalt driving lanes show damage due to subsidence (depressions of a scale >50 cm) or heavy traffic (shallow ruts in asphalt caused by trucks in summer). This means that the road can be used by normal city bikes, wheelchairs and sports cars, but not by a racing bike.
bad Damaged asphalt roads that show clear signs of maintenance: This might include potholes, some of them quite deep, which might decrease the average speed of cars. However, it isn’t so rough that ground clearance becomes a problem. Meaning that the street causes trouble to normal city bikes but not a trekking bike and a car.
A.2 Concrete
Quality Description
excellent As good as new concrete, on which a skateboard or rollerblades will have no problem.
good Concrete road showing the first signs of wear, such as narrow, smaller than 1.5 cm cracks, or wider cracks filled up with tar, shallow dents in which rainwater may collect, which may cause trouble for rollerblades but not for racing bikes.
intermediate Concrete roads that show signs of maintenance, such as patches of repaired surface, wider cracks larger than 2cm. Concrete sidewalks may contain potholes, but these are small, shallow (<3cm deep) and can be easily avoided. Concrete driving lanes show damage due to subsidence (depressions of a scale >50 cm) or heavy traffic (shallow ruts in concrete caused by trucks in summer). This means that the road can be used by normal city bikes, wheelchairs and sports cars, but not by a racing bike.
bad Heavily damaged concrete roads that badly need maintenance: many potholes, some of them quite deep. The average speed of cars is less than 50% of what it would be on a smooth road. However, it isn’t so rough that ground clearance becomes a problem. Meaning that the street causes trouble to normal city bikes but not a trekking bike and a car.
A.3 Paving stones
Quality Description
excellent Newly installed and regularly laid paving stones that show no signs of wear. Gaps may be visible, but they are small and uniform and do not significantly affect the driving experience.
good Paving stones showing first signs of wear or newly installed stones with visible but uniform gaps between them. While still suitable for most activities, these surfaces may pose minor challenges for rollerblades and skateboards but remain navigable for racing bikes.
intermediate Characterized by paving stones exhibiting multiple signs of wear, such as shifted heights, potholes, or cracks. This grade allows for the comfortable passage of normal city bikes and standard vehicles but may prove challenging for racing bikes.
bad Heavily uneven or damaged paving stones in dire need of maintenance, featuring significant height disparities and numerous deep potholes. While ground clearance remains sufficient for most vehicles, the surface severely impedes travel, particularly for standard city bikes.
A.4 Sett
Quality Description
good The best sett roads with flattened stones and gaps that are at most small or filled up. The surface might cause trouble for rollerblades but not for racing bikes.
intermediate The surfaces of the sett stones are not completely flat, or there may be slightly larger gaps between the stones, this causes problems for racing bikes and slows down city bikes and cars.
bad Sett stones with large and possibly uneven gaps or uneven stone surfaces or damaged stones, resulting in an overall bumpy surface: This results in a highly uncomfortable driving experience for city bikes. The average speed of cars is less than 50% of what it would be on a smooth road. However, it isn’t so rough that ground clearance becomes a problem.
A.5 Unpaved
Quality Description
intermediate The best unpaved roads that have a compacted surface. This grade allows for the passage of normal city bikes and standard vehicles but may prove challenging for racing bikes.
bad Unpaved roads that do not have a smooth and compacted surface but ones that can still be used with a car or trekking bike but not with a city bike. This category also includes hiking paths, which are too narrow for cars.
very_bad Unpaved roads with potholes, ruts or generally a highly uneven surface not safely passable with a regular passenger car but still passable with an average SUV with higher ground clearance. This category also includes highly uneven hiking paths, which are too narrow for cars.
Appendix B GPT 4-o Prompt
The following shows the utilized prompt for one-shot asphalt image prediction using GPT 4-o. Prompts for other surface types are adjusted according to type-specific labeling schemes (see Appendix  A).
    "role": "system",
    "content": [
            "type": "text",
            "text": "You are a data annotation expert trained to
            classify the quality level of road surfaces in images."
    "role": "user",
    "content": [
            "type": "text",
            "text": f"""
                    You need to determine the quality level of the road
                    surface depicted in the image, following this defined scale:
                    Asphalt surfaces are graded from excellent to bad
                    according to the following scale:
                    1) excellent: As good as new asphalt, on which a
                    skateboard or rollerblades will have no problem.
                    2) good: Asphalt showing the first signs of wear, such as
                    narrow, smaller than 1.5 cm cracks, or wider cracks filled
                    up with tar, shallow dents in which rainwater may collect,
                    which may cause trouble for rollerblades but not for
                    racing bikes.
                    3) intermediate: Asphalt roads that shows signs of
                    maintenance, such as patches of repaired surface, wider
                    cracks larger than 2cm. Asphalt sidewalks may contain
                    potholes, but these are small, shallow (<3cm deep) and can
                    be easily avoided, asphalt driving lanes shows damage due
                    to subsidence (depressions of a scale >50 cm) or heavy
                    traffic (shallow ruts in asphalt caused by trucks in
                    summer). This means that the road can be used by normal
                    city bikes, wheelchairs and sports cars, but not by a
                    racing bike.
                    4) bad: Damaged asphalt roads that show clear signs of
                    maintenance: This might include potholes, some of them
                    quite deep, which might decrease the average speed of
                    cars.  However, it isnt so rough that ground clearance
                    becomes a problem. Meaning that the street causes trouble
                    to normal city bike but not a trekking bike and a car.
                    Please adhere to the following instructions:
                    1) Step 1: If you detect multiple surface types, only
                    consider the path, driving lane, cycleway or sidewalk in
                    the focus area.
                    2) Step 2: Check if the road surface is worn off and if
                    you can find any damages, like cracks.
                    3) Step 3: Check the quantity and the size of the damages.
                    3) Step 4: Then decide if you could ride on the surface
                    with  a skateboard, rollerblades, racing bikes, city bike,
                    or a normal car.
                    4) Step 5: If you detect characteristics of two classes,
                    choose the worse class.
                    How would you rate this image using one of the four
                    options of the defined scale:
                    1) excellent
                    2) good
                    3) intermediate
                    4) bad
                    Provide your rating in one word disregarding the bullet
                    point numbers and brackets as a string using the four
                    levels of the scale provided.
                    Make sure you have the same number of image urls as input
                    as you have output values.
                    Do not provide any additional explanations for your
                    rating; focus solely on the road surface quality."""
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{excellent_encoded_image}"
            "type": "text",
            "text": "This was an example for ’excellent’"
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{good_encoded_image}"
            "type": "text",
            "text": "This was an example for ’good’"
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{intermediate_encoded_image}"
            "type": "text",
            "text": "This was an example for ’intermediate’"
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{bad_encoded_image}"
            "type": "text",
            "text": "This was an example for ’bad’"
            "type": "text",
            "text": "Please decide now in one word which category the
            following picture belongs to as instructed in the beginning of the
            prompt. Then compare this image to the previous ones and decide
            whether this category is correct."
            "type": "image_url",
            "image_url": {
                        "url": f"data:image/jpeg;base64,{image}"

### v3

@@ Methods
Generally, performance increases of GPT-4o and similar models are to be expected in the future, further amplifying the viability of this approach. To reduce dependency on OpenAI and monetary cost, future work should evaluate open-source alternatives. Similarity search is more efficient in computational and monetary terms, therefore currently remaining a viable alternative despite inferior results. The results varied between classes and showed the worst performance for paving stones - bad, where there was no improvement to the baseline. Further experiments, especially with more images for the smallest class, could be explored, as well as advancements such as incorporating clustering strategies [32].
Data Records
StreetSurfaceVis is an image dataset containing 9,122 street-level images within Germany’s bounding box with labels on road surface type and quality; find the number of instances per class in Table 4. A csv file contains all the image metadata, and four folders contain the image files. Based on the image width, all images are available in four different sizes: 256px, 1024px, 2048px, and the original size. Folders containing the images are named according to the respective image size. Image files are named based on the mapillary_image_id. This repository (https://doi.org/10.5281/zenodo.11449977) provides the dataset, a description, the labeling guide, and a datasheet documenting the dataset [33].
Technical Validation
Inter-rater reliability
To evaluate inter-rater reliability, 180 images (10 images per class according to OSM pre-labels) were independently rated by all three annotators. Twelve images were marked for revision by at least one annotator, and another 49 were discarded by at least one annotator for the above reasons, all of which were excluded from the calculation of inter-rater reliability. Krippendorff’s $\alpha$ [34] for surface type is calculated at 0.96, indicating a high level of agreement. Surface quality, treated as an ordinal scale variable, achieves a Krippendorff’s $\alpha$ of 0.74. While this is generally deemed an acceptable level of agreement, it reflects the fluid class transitions of quality in contrast to type.
Type and quality model performance
To assess the validity of our dataset, we train an EfficientNetV2-S-based model to predict surface types. We split the final dataset into a training set of 8,346 images and a test set of 776 images from five cities geographically distinct from the training data. Note that we do not enhance underrepresented classes in the test data, aiming to reflect real-life distributions. traapplyvan 80:20lidation split of 80:20 and conduct five runs with different seeds, especially influencing the train-validation split, and report the averaged results.
We use the validation dataset solely to identify the optimal number of epochs, without tuning other hyperparameters. An accuracy (loss) of 0.96 (0.13) is achieved for the training data, 0.94 (0.19) for validation, and 0.91 for test data, respectively. Table 6 presents the recall, precision, and F1 scores of the test data for each surface type. All F1 scores for the test data are equal to or exceed 0.9, except for the ‘concrete’ surface type. This demonstrates a strong generalization of our training data to Mapillary images from previously unseen (German) cities. The low F1 score of 0.35 for concrete can be attributed to its visual similarity to asphalt and its rare occurrence in the dataset. Consequently, a small portion of the large asphalt class is misclassified as concrete. Given the limited number of concrete images, this results in a low precision for the concrete class. Depending on the application, e.g., surface classification for routing purposes, distinguishing between concrete and asphalt may not be needed, and the two classes could be merged into one.
In a similar setup, we train five regression models, one for each type, to predict surface quality, using mean squared error as the loss function. For evaluation, we assume a correctly classified type, i.e., the quality prediction is independent of the type model. Deviations from the true value are normally distributed and centered around 0, with an overall Spearman correlation coefficient of 0.72 and type-specific coefficients between 0.42 and 0.65 (see Table 6), thus moderate to strong correlations. When converting numeric predictions into quality categories, the models achieve an overall accuracy of 0.63, with type-specific accuracies ranging from 0.61 to 0.71. To account for fluidity in quality annotation, we also compute the 1-off accuracy which considers neighboring classes as correct classifications. All 1-off accuracies are (almost) 1.0, showing that all model predictions are at most one class off, demonstrating a high level of precision.
Cross-dataset generalization testing
To demonstrate the ability of our training dataset to train models that generalize to other data sources, we train the type and quality models on our dataset to predict the RTK [14] dataset and vice versa. RTK contains low-resolution street-level images captured with one moving vehicle in a Brazilian town. Images are labeled according to type (asphalt, paved, and unpaved) and quality (good, regular and bad), resulting in the following instances: asphalt-good: 1,978, asphalt-regular: 839, asphalt-bad: 464, paved-good: 1,179, paved-regular: 324, paved-bad: 124, unpaved-bad: 593, and unpaved-regular: 796.
We merge our asphalt and concrete images to a single class, matching the RTK ‘asphalt’ class, and, similarly, paving stones and sett are merged into ‘paved’. As we formulate the quality prediction as a regression problem, we can utilize the Spearman correlation coefficient to compare true and predicted values and thus do not need to match labels. Since our dataset is larger, we down-sample our dataset to match the RTK image count of 6,297 when utilized for training while maintaining our class distribution. Again, we train each model five times with different seeds and report averaged results. As there are different class sizes between models, we report the average (unweighted) F1 score as an overall metric. We determine significance according to a two-sided Mann-Whitney U test (nonparametric alternative of the t-test) with a significance level of 0.05.
As shown in Table 7, the model trained on our dataset achieves a significantly higher average F1 score of 0.81 compared to 0.56 of the model trained on RTK. While the RTK model slightly outperforms our dataset on the recall of paved roads, it performs poorly on the detection of unpaved roads. Note that we trained the models in a vanilla setting, without applying additional techniques such as blurring augmentation, which is expected to enhance the performance of the model trained on StreetSurfaceVis due to the higher resolution of the images.
To compare surface quality predictions, we consider the Spearman correlation coefficient between true types and model predictions: overall model predictions based on our dataset achieve a coefficient of 0.52, while it is 0.16 vice versa, indicating that the model based on our dataset captures quality differences of the RTK dataset, while not the other way round. See Table 7 for type-wise correlation coefficients. Note, that quality prediction works best for asphalt roads which is not surprising as the quality definitions for paved and unpaved do not entirely align between both datasets.
Usage Notes
Train-test split
For modeling, we recommend using a train-test split where the test data includes geospatially distinct areas, thereby ensuring the model’s ability to generalize to unseen regions is tested. We propose urban areas of five cities varying in population size and from different regions in Germany for testing, comprising 776 images tagged accordingly.
Cropping
As the focal road located in the bottom center of the street-level image is labeled, we recommend to crop images to their lower and middle half prior using for classification tasks.
This is an exemplary code for recommended image preprocessing in Python:
    from PIL import Image
    img = Image.open(image_path)
    width, height = img.size
    img_cropped = img.crop((0.25 * width, 0.5 * height, 0.75 * width, height))
License
Images are obtained from Mapillary, a crowd-sourcing plattform for street-level imagery. More metadata about each image can be obtained via the Mapillary API . User-generated images are shared by Mapillary under the CC-BY-SA License. For each image, the dataset contains the mapillary_image_id and user_name. User information can be accessed on the Mapillary website by https://www.mapillary.com/app/user/<USER_NAME> and image information by https://www.mapillary.com/app/?focus=photo&pKey=<MAPILLARY_IMAGE_ID>. To use the provided images, it is required to adhere to the terms of use of Mapillary [^7] https://www.mapillary.com/terms?locale=de_DE.
Code availability
The code for image selection, download and preparation for manual annotation is provided in this repository: https://github.com/SurfaceAI/dataset_creation.